ECH 4323: Chemical Process Control SIMULATION PROJECT

Propylene glycol is produced by the hydrolysis of propylene oxide in a CSTR. The reaction is normally conducted using excess water. To get propylene glycol of the desired final specification, the reaction products are sent through a dewatering distillation column.

The objective of this project is to analyze the steady-state and dynamic behavior of the propylene glycol process and design suitable controllers for the reactor as well as the distillation column.

The following data is available for the reactor part of the process:

1. Reactor volume: 1 m^{3}

2. Feed Characteristics:

• Propylene oxide: F_{A0} kmol/hr

• Methanol: 33 kmol/hr

• Water (with 0.1 wt. % H_{2}SO_{4}): 365 kmol/hr

• Feed Temperature: T_{0} ^{o}C

3. The reaction is first order with respect to propylene oxide

−r_{A} = kC_{A}

where

\begin{aligned}k & = & Aexp (−\frac{E}{RT} ) \\ A & = & 16.96 × 10^{12}hr^{−1} \\ E & = & 75, 200 kJ/kmol \end{aligned}

The heat of reaction as a function of temperature is given by

ΔH_{rxn} = −84, 500 − 29.24(T − 293) kJ/kmol propylene oxide

where T is the reactor temperature in K.

4. The temperature cannot exceed 52^{o}C to prevent excessive loss of propylene oxide due to evaporation. For this reason, there is a cooling jacket (with water as the coolant) around the reactor with the following characteristics:

(a) Coolant inlet temperature: 15^{o}C

(b) Coolant flow rate: ṁ_{co}

(c) (Overall Heat transfer coeff.)(Area): UA = 3000 kJ/kgmol/K

Additional data on reactants and products: Density:

Propylene Oxide 14.95 kmol/m^{3}
Methanol  24.71 kgmol/m^{3}
Water  55.37 kgmol/m^{3}

Specific Heat:

Propylene Oxide 146 kJ/kmol/K
Methanol  81 kJ/kmol/K
Water  75 kJ/kmol/K
Propylene Glycol 192 kJ/kmol/K

1. Compute the steady state values of propylene oxide, propylene glycol, and the reactor temperature for the following conditions (you may use MATLAB or hand calculations for this part):

Coolant flow, ṁ_{co} Feed Temp., T_{0} P.O. in Feed, F_{A0}
kmol/hr  ^{o}C  kmol/hr
450  24  20
450  20  20
450  28  20
400  24  20
500  24  20
450  24  16
450  24  24

2. Suppose at time t = 0, the reactor is at the same conditions as the feed characteristics of the first line of the above table. Develop a dynamic model for the reactor system. Solve for the time profiles of propylene oxide, propylene glycol and the reactor temperature as a function of time in MATLAB.

3. Suppose the inlet temperature undergoes the following pulse change:

T_{0} = \{ \begin{matrix}24 & 0 ≤ t < 2.5 hr \\ 28 & 2.5 ≤ t < 3 hr \\ 24 & t ≥ 3 hr \end{matrix}  (1)

How do the reactor temperature and the concentrations of propylene oxide and propylene glycol change with time? Show the appropriate plots in MATLAB.

4. Is the CSTR stable under the following nominal conditions?

Coolant flow  450 kmol/hr
Feed Temperature  24 ^{o}C
Propylene Oxide in Feed 20 kmol/hr

Perform a linear stability analysis to answer the above question.

5. Design a distillation column in CHEMCAD that takes the reactor outlet under nominal conditions and gives a bottoms product that is 99.5% propylene glycol. Show the CHEMCAD simulation results.

6. What happens in the CHEMCAD simulation when the propylene glycol in the feed to the column is increased or decreased by 10%?

7. Suppose the CSTR is operating at a coolant flow rate of 450 kmol/hr and a propylene oxide feed rate of 20 kmol/hr. The initial feed temperature is 20 ^{o}C. Design a PI controller in MATLAB that keeps the outlet proplyne glycol concentration at 0.50 kmol/m^{3}. Show the plots of the feed temperature versus time and the proplyne glycol concentration in the outlet versus time.

8. Implement the above controller in CHEMCAD.